OpenStudy (anonymous):

A standard number cube is tossed. Find P(2 or even). plz help

OpenStudy (anonymous):

P(2)=1/6(as we have six faces and one dot means choosing 1 out of 6) P(even)=1/2 (as there are 3 even out of 6) P(2 or even)=P(2)+P(even)-P(2 and even) =1/6+1/3-1/6=1/3

OpenStudy (anonymous):

thank you